What is the output of the following code? for(int i =0; i<3; i++) { int x = 0; x = x+i; Sy System.out.println(x);
Answer #1

Here, we have given for loop which intially takes value of i = 0 then it checks condition i<3. As the condition is true (0 < 3) , so loop executes the statement in it. The first statement takes x = 0 then second statement will add up the value of x to i and then final value of x becomes 0. Now i will increase by 1 and will become 1 as 0 + 1 = 1. Now condition will be checked as earlier for i<3 which gives 1<3. Since the condition is true so the loop executes again which provides x = 0 and then adding x and i which will make x = 0 + 1 = 1. Now i will again increase by 1 and will become 1 + 1 = 2. After this condition will be checked as i < 3 that is 2<3. As the condition is true, the loop executes and will give x = 0 and adding x and i to give x = 2 + 0=2. Now i will increase again and become 3(2 + 1 = 3). When condition will be checked that is 3 < 3, it becomes false. So loop will not be executed further and the control will come out of the loop to next statement after the loop which prints the value of x = 2 which is the stored value of x variable during the last time of execution of loop. So in this way the for loop executes in 3 steps: Declaration of variable, Condition Check and Increment or Decrement of variable.